### v1

# Open cell property in weakly o-minimal structures

Abstract.
In an o-minimal structure, Wilkie proved that every bounded definable open set is a finite union of definable open cells. We generalze it to weakly o-minimal structures.
Key words and phrases:
Weakly o-minimal structures, open strong cells and open cell property
2020 Mathematics Subject Classification:
Primary 03C64
1. Introduction
Let $\mathcal{M}=(M,<,\ldots)$ be an expansion of a dense linear order without endpoints. A subset $A$ of $M$ is said to be convex if $a,b\in A$ and $c\in M$ with $a<c<b$ then $c\in A$. Moreover if $A=\emptyset$ or $\inf A,{\rm sup}A\in M\cup\{-\infty,+\infty\}$, then $A$ is called an interval in $M$. We say that $\mathcal{M}$ is o-minimal (weakly o-minimal) if every definable subset of $M$ is a finite union of intervals (convex sets), respectively.
In o-minimal structures, Wilkie ([9]) proved every bounded definable open set is a finite union of open cells. Andrews ([1]) proved that if an o-minimal structure admits CE-cell decomposition then any definable open set is expressed as a finite union of definable open cells. Edmundo et al ([3]) proved in semi-bounded o-minimal expansion of an ordered group, any definable open set is covered a finite union of open cells. The relation between strong cell decomposition in weakly o-minimal structures and open cell property in o-minimal structures is proved by [4]. We consider a generalization of the above results in weakly o-minimal structures.
2. Weakly o-minimal expansions of real closed fields
Let $\mathcal{M}=(M,<,\ldots)$ be an weakly o-minimal structure. For any subsets $C,D$ of $M$, we write $C<D$ if $c<d$ whenever $c\in C$ and $d\in D$. A pair $\langle C,D\rangle$ of non-empty subsets of $M$ is called a cut in $\mathcal{M}$ if $C<D,C\cup D=M$ and $D$ has no lowest element. A cut $\langle C,D\rangle$ is said to be definable in $\mathcal{M}$ if the sets $C,D$ are definable in $\mathcal{M}$. The set of all cuts definable in $\mathcal{M}$ will be denoted by $\overline{M}$. Note that we have $M=\overline{M}$ if $\mathcal{M}$ is o-minimal. We define a linear ordering on $\overline{M}$ by $\langle C_{1},D_{1}\rangle<\langle C_{2},D_{2}\rangle$ if and only if $C_{1}\subsetneq C_{2}$. Then we may treat $(M,<)$ as a substructure of $(\overline{M},<)$ by identifying an element $a\in M$ with the definable cut $\langle(-\infty,a],(a,+\infty)\rangle$. Morever, if $\mathcal{M}=(M,<,+,\ldots)$ is an expansion of an ordered group $(M,<,+)$, then the cut $\langle C,D\rangle$ is said to be valuational if there exists $\varepsilon>0$ such that for all $x\in C$ and $y\in D$, we have $y-x>\varepsilon$. Otherwise the cut $\langle C,D\rangle$ is said to be non-valuational. The structure $\mathcal{M}$ is said to be non-valuational if all definable cuts in $\mathcal{M}$ are non-valuational.
We equip $M$ ($\overline{M}$) with the interval topology (the open intervals form a base), and each product $M^{n}$ ($(\overline{M})^{n}$) with the corresponding product topology, respectively.
Recall the notion of definable functions from [8]. Let $n$ be a positive integer and $A\subseteq M^{n}$ definable. A function $f:A\to\overline{M}$ is said to be definable if the set $\{\langle x,y\rangle\in M^{n+1}:x\in A,y<f(x)\}$ is definable. A function $f:A\to\overline{M}\cup\{-\infty,+\infty\}$ is said to be definable if $f$ is a definable function from $A$ to $\overline{M}$, $f(x)=$-$\infty$ for all $x\in A$, or $f(x)=+\infty$ for all $x\in A$. A definable function $f:A\to\overline{M}\cup\{-\infty,+\infty\}$ is said to be regular if it takes all its values in one of the sets $M$, $\overline{M}-M$, $\{-\infty\}$, or $\{+\infty\}$.
Definition 2.1 ([7]).
For every $m\in\mathbb{N}_{+}$, we define, by induction, strong cells $C\subseteq M^{m}$ and their completions $\overline{C}\subseteq\overline{M}^{m}$.
  1. (1)
    Any singleton of $M$ is a 0-strong cell in $M$ and is equal to its completion.
  2. (2)
    Any non-empty open convex definable subset of $M$ is a 1-strong cell in $M$. If $C\subseteq M$ is a 1-strong cell, then we define its completion by $\overline{C}:=\{x\in\overline{M}:(\exists a,b\in C)(a<x<b)\}$.
    Let $m\in\mathbb{N}_{+}$, $k\leq m$ and suppose that we have already defined $k$-strong cells in $M^{m}$ and their completions in $\overline{M}^{m}$.
  3. (3)
    If $C\subseteq M^{m}$ is a $k$-strong cell and $f:C\to M$ is a continuous definable function which has a unique continuous extension $\overline{f}:\overline{C}\to\overline{M}$, then the graph of $f$ which is denoted by $\Gamma(f)_{C}$, is a $k$-strong cell in $M^{m+1}$ and its completion in $\overline{M}^{m+1}$ is defined as $\Gamma(\overline{f})_{\overline{C}}$.
  4. (4)
    If $C\subseteq M^{m}$ is a $k$-strong cell and $f,g:C\to M\cup\{\pm\infty\}$ is continuous definable functions such that $f,g$ have continuous extensions $\overline{f},\overline{g}:\overline{C}\to\overline{M}$, where $\overline{f}(\overline{a})<\overline{g}(\overline{a})$ for any $\overline{a}\in\overline{C}$, then the set $(f,g)_{C}:=\{\langle\overline{a},b\rangle\in C\times M:f(\overline{a})<b<g(\overline{a})\}$ is called a $(k+1)$-strong cell in $M^{m+1}$. The completion of $(f,g)_{C}$ in $\overline{M}^{m+1}$ is defined as $(\overline{f},\overline{g})_{\overline{C}}:=\{\langle\overline{a},b\rangle\in\overline{C}\times\overline{M}:\overline{f}(\overline{a})<b<\overline{g}(\overline{a})\}$.
A strong cell $C\subseteq M^{m}$, $m\geq 2$, is called a refined strong cell if each of the boundary functions appearing in its definition is regular. Refined strong cells in $M$ coincide with strong cells in $M$.
Let $C\subseteq M^{m}$ be a strong cell and $f:C\to\overline{M}$ definable. The function $f$ is said to be strongly continuous if it has a unique continuous extension $\overline{f}:\overline{C}\to\overline{M}$.
Definition 2.2.
Any finite partition of $M$ into singletons and convex open sets definable in $\mathcal{M}$ is called a strong cell decomposition of $M$. A finite partition $\mathcal{C}$ of $M^{m+1}$ into strong cells is said to be a strong cell decomposition of $M^{m+1}$ if $\pi[\mathcal{C}]:=\{\pi[C]:C\in\mathcal{C}\}$ is a strong cell decomposition of $M^{m}$. Here $\pi:M^{m+1}\to M^{m}$denotes the projection dropping the last coordinate. A strong cell decomposition $\mathcal{C}$ of $M^{m}$ partitions $X\subseteq M^{m}$ if for any $C\in\mathcal{C}$, we have $C\subseteq X$ or $C\cap X=\emptyset$. The structure $\mathcal{M}$ has the strong cell decomposition property if for any positive integers $m$, $k$ and any definable sets $X_{1},\ldots,X_{k}\subseteq M^{m}$, there exists a decomposition of $M^{m}$ into finitely many strong cells partitioning each of the sets $X_{1},\ldots,X_{k}$. Similarly, we can define the refined strong cell decomposition property.
The following result follows from Lemma 2.6 and Corollary 2.16 of [7].
Fact 2.3.
Let $\mathcal{M}=(M,<,+,\ldots)$ be a weakly o-minimal expansion of an ordered group. Then the structure $\mathcal{M}$ is a non-valuational if and only if it has a strong cell decomposition property.
Let $\mathcal{C}$ and $\mathcal{D}$ are refined strong cell decompositions of $M^{m}$. We say that $\mathcal{D}$ refines $\mathcal{C}$ (denoted by $\mathcal{C}\prec\mathcal{D}$) if every cell in $\mathcal{C}$ is a union of some cells in $\mathcal{D}$
The following statement is Fact 2.1 of [8].
Fact 2.4.
Let $k,m\in\mathbb{N}_{+}$. If $X_{1},\ldots,X_{k}\subseteq M^{m}$ are definable sets, then there exists a smallest refined strong cell decomposition $\mathcal{D}$ of $M^{m}$ partitioning each of $X_{1},\ldots,X_{k}$.
For any $n\in\mathbb{N}_{+}$ and any refined strong cell $C\subseteq M^{n}$, we denote by $R_{C}$ an $n$-ary relational symbol. We interpret $R_{C}$ in $\overline{M}^{n}$ as $\overline{C}$, the completion of $C$. According to Section 2 of [8], the structure $\overline{\mathcal{M}}:=(\overline{M},<,(R_{C}:C\text{ is a refined strong cell}))$ is o-minimal and is called the canonical o-minimal extension of $\mathcal{M}$. If $X\subseteq\overline{M}^{m}$ is a set definable in $\overline{\mathcal{M}}$, then $X\cap M^{m}$ is definable in $\mathcal{M}$. If additionally $Y\subseteq M^{m}$ is definable in $\mathcal{M}$, then $X\cap Y$ is definable in $\mathcal{M}$.
The following definition is Definition 3.1 of [6].
Definition 2.5.
Let $m\in\mathbb{N}_{+}$. Suppose that $X$ is a defianble subset of $M^{m}$ and $\mathcal{C}$ is the smallest refined strong cell decomposition of $M^{m}$ partitioning $X$. We define the completion of $X$ in $\overline{M}^{m}$ as $\overline{X}:=\bigcup\{\overline{C}:C\in\mathcal{C}\wedge C\subseteq X\}$.
Remark 2.6.
Note that there exist sets $C,D$ such that $C\subseteq D$ but $\overline{C}\not\subseteq\overline{D}$. Let $R_{alg}$ denote the ordered field of real algebraic numbers and let the unary predicate symbol $P$ be interpreted by the convex set $S=(-\pi,\pi)\cap R_{alg}$. Then, by Proposition 2.1 from [5], the structure $\mathfrak{R}=(R_{alg},<,+,\cdot,P)$ is non-valuational weakly o-minimal. So the structure $\mathfrak{R}$ has the refined strong cell decomposition property. Consider the functions $f,g:I=(0,4)\to\mathfrak{R}$ defined by $f(x)=x$ and $g(x)=x+1$, and the function $h:J=(0,\pi)\to\mathfrak{R}$ defined by $h(x)=x+2$. Let $C=\Gamma(f)_{I}\cup\Gamma(g)_{I}$ and $D=\Gamma(f)_{I}\cup\Gamma(g)_{I}\cup\Gamma(h)_{J}$. Then, we have $C\subseteq D$ but $\overline{C}\not\subseteq\overline{D}$.
Remark 2.7.
Let $m\in\mathbb{N}_{+}$. Suppose that $X$ is a defianble subset of $M^{m}$. Let $\mathcal{C}$ be the smallest refined strong cell decomposition of $M^{m}$ partitioning $X$ such that $X=\bigcup_{i=1}^{k}C_{i}$ and $\overline{X}=\bigcup_{i=1}^{k}\overline{C_{i}}$ with $C_{1},\ldots,C_{k}\in\mathcal{C}$. If $X$ is open in $M$, then $\overline{X}$ is open in $\overline{M}$.
Proposition 2.8 (Weak definable choice).
Suppose that $\mathcal{M}=(M,<,+,\ldots)$ is a weakly o-minimal expansion of an ordered group $(M,<,+)$. Let $S\subseteq M^{m+n}$ be definable and $\pi:M^{m+n}\to M^{m}$ the projection of the first $m$ coordinates. Then, there exists a definable map $f:\pi(S)\to\overline{M}^{n}$ such that $\Gamma(f)\subseteq\overline{S}$.
Proof.
Let $\mathcal{C}$ be the smallest refined strong cell decomposition of $M^{m+n}$ partitioning $S$ such that $S=\bigcup_{i=1}^{k}C_{i}$ and $\overline{S}=\bigcup_{i=1}^{k}\overline{C_{i}}$ with $C_{1},\ldots,C_{k}\in\mathcal{C}$. By the o-minimality of $\overline{{\mathcal{M}}}$ and Proposition 6.1.2 of [2], for any $C\in\mathcal{C}$ with $C\subseteq S$, there exists a map $\overline{f}_{C}:\overline{\pi(C)}\to\overline{M}^{n}$ such that $\Gamma(\overline{f}_{C})\subseteq\overline{C}$ and $\overline{f}_{C}$ is definable in $\overline{{\mathcal{M}}}$. By Propisition 2.3 from [8], $\overline{f}_{C}|_{\overline{\pi(C)}\cap M^{m}}:\overline{\pi(C)}\cap R^{m}\to\overline{M}^{n}$ is definable in $\mathcal{M}$}. By $S=\bigcup_{i=1}^{k}C_{i}$ and $\overline{S}=\bigcup_{i=1}^{k}\overline{C_{i}}$, this completes the proof. ∎
Proposition 2.9.
Suppose that $\mathcal{M}=(M,<,+,\ldots)$ is a weakly o-minimal expansion of an ordered group $(M,<,+)$. Let $S\subseteq M^{m+n}$ be definable and $\overline{\pi}:\overline{M}^{m+n}\to\overline{M}^{m}$ the projection of the first $m$ coordinates. Then, there exists a map $\overline{g}:\overline{\pi}(\overline{S})\to\overline{S}$ such that $\overline{\pi}\circ\overline{g}=\mathrm{id}_{\overline{\pi}(\overline{S})}$ and $\overline{g}$ is definable in $\overline{{\mathcal{M}}}$.
Proof.
Let $\pi:=\overline{\pi}|_{M^{m+n}}$. Suppose that $\mathcal{C}$ is the smallest refined strong cell decomposition of $M^{m+n}$ partitioning $S$ such that $S=\bigcup_{i=1}^{k}C_{i}$ and $\overline{S}=\bigcup_{i=1}^{k}\overline{C_{i}}$ with $C_{1},\ldots,C_{k}\in\mathcal{C}$. By the o-minimality of $\overline{{\mathcal{M}}}$ and Proposition 6.1.2 of [2], for any $C\in\mathcal{C}$ with $C\subseteq S$, there exists a map $\overline{f}_{C}:\overline{\pi(C)}\to\overline{M}^{n}$ such that $\Gamma(\overline{f}_{C})\subseteq\overline{C}$ and $\overline{f}_{C}$ is definable in $\overline{{\mathcal{M}}}$. Let $\overline{f}=\bigcup_{i=1}^{k}\overline{f}_{C_{i}}$. Then we have $\overline{f}:\overline{\pi}(\overline{S})\to\overline{M}^{n}$ and $\Gamma(\overline{f})\subseteq\overline{S}$. We define the $\overline{\mathcal{M}}$-definable map $\overline{g}:\overline{\pi}(\overline{S})\to\overline{S}$ by $\overline{g}(x)=(x,\overline{f}(x))$. Then it follows $\overline{\pi}\circ\overline{g}=\mathrm{id}_{\overline{\pi}(\overline{S})}$. ∎
Example 2.10.
Let $R_{alg}$ denote the ordered field of real algebraic numbers and let the unary predicate symbol $P$ be interpreted by the convex set $S=(-\pi,\pi)\cap R_{alg}$. Then, by Proposition 2.1 from [5], the structure $\mathfrak{R}=(R_{alg},<,+,\cdot,P)$ is non-valuational weakly o-minimal. So the structure $\mathfrak{R}$ has the refined strong cell decomposition property. Consider the functions $f,g:I=(-1,1)\to\overline{\mathfrak{R}}$ defined by f(x)=4f(x)=-4 and $g(x)=\pi x$. Then, the set $(f,g)_{I}$ is an open strong cell but not an open refined strong cell. Moreover, $(f,g)_{I}$ cannot be expressed as a finite union of open refined strong cells.
Lemma 2.11 ([5]).
Let $\mathcal{M}$ be a weakly o-minimal expansion of an ordered field. Then $\mathcal{M}$ is a real closed field.
Fact 2.12 ([9]).
Suppose that $\mathcal{M}=(M,<,+,\cdot,\ldots)$ is an o-minimal expansion of a real closed field. Let $U$ be a definable bounded open subset of $M^{n}$. Then, there exists a finite collection of open cells in $M^{n}$ whose union is $U$.
Theorem 2.13.
Suppose that $\mathcal{M}=(M,<,+,\cdot,\ldots)$ is a non-valuational weakly o-minimal expansion of a real closed field. Let $U$ be a definable bounded open subset of $M^{n}$. Then, there exists a finite collection of open strong cells in $M^{n}$ whose union is $U$.
Proof.
Let $\mathcal{C}$ be the smallest refined strong cell decomposition of $M^{n}$ partitioning $U$ such that $U=\bigcup_{i=1}^{k}C_{i}$ and $\overline{U}=\bigcup_{i=1}^{k}\overline{C_{i}}$ with $C_{1},\ldots,C_{k}\in\mathcal{C}$. By Remark 2.7, the set $\overline{U}$ is a definable bounded open subset of $\overline{M}^{n}$. Then, by Fact 2.12, there exist open cells $D_{1},\dots,D_{l}$ such that $\overline{U}=D_{1}\cup\dots\cup D_{l}$. Since each $D_{i}\cap M^{n}$ is an open strong cell and ${U}=(D_{1}\cap M^{n})\cup\dots\cup(D_{l}\cap M^{n})$, this completes the proof. ∎
3. Weakly o-minimal expansions of ordered groups
Let $\mathcal{M}$ be a non-vaulational weakly o-minimal expansion $\mathcal{M}=(M,<,+,0,\cdots)$ of an ordered group. We say that $\mathcal{M}$ is semisemi-boundedbounded if there exists no definable bijection between a bounded interval and a unbounded interval.
Theorem 3.1 ([3]).
Let $\mathcal{M}$ be a semi-bounded o-minimal expansion $\mathcal{M}=(M,<,+,0,\cdots)$ of an ordered group. Then every non-empty definable open set is a finite union of open cells.
Theorem 3.2.
Consider a semi-bounded non-valuational weakly o-minimal expansion $\mathcal{M}=(M,<,+,\dots)$ of an ordered group. Let $U$ be a definable open subset of $M^{n}$. Then there exist finitely many strong open cells whose union is $U$.
Proof.
Take the smallest refined strong cell decomposition $\mathcal{C}$ of $M^{n}$ partitioning $U$ such that $U=\bigcup_{i=1}^{k}C_{i}$ and $\overline{U}=\bigcup_{i=1}^{k}\overline{C_{i}}$ with $C_{1},\ldots,C_{k}\in\mathcal{C}$. By Remark 2.7, the set $\overline{U}$ is a definable open subset of $\overline{M}^{n}$. Then, by Theorem 3.1, there exist open cells $D_{1},\dots,D_{l}$ such that $\overline{U}=D_{1}\cup\dots\cup D_{l}$. Since each $D_{i}\cap M^{n}$ is an open strong cell and ${U}=(D_{1}\cap M^{n})\cup\dots\cup(D_{l}\cap M^{n})$, this completes the proof. ∎

### v2

@@
 Abstract.
-In an o-minimal structure, Wilkie proved that every bounded definable open set is a finite union of definable open cells. We generalze it to weakly o-minimal structures.
+Every bounded definable open set is a union of finitely many open strong cells in a weakly o-minimal expansion of a real closed field. We prove this fact and another theorem similar to it.
 Key words and phrases:
-Weakly o-minimal structures, open strong cells and open cell property
+Weakly o-minimal structures, open strong cells and open cell property. The first author was partially supported by JSPS KAKENHI Grant Number 23K03095.
 2020 Mathematics Subject Classification:
@@
 1. Introduction
-Let $\mathcal{M}=(M,<,\ldots)$ be an expansion of a dense linear order without endpoints.
+Throughout this paper, “definable” means “definable possibly with parameters” and we assume that a structure $\mathcal{M}=(M,<,\dots)$ is a dense linear order without endpoints.
 A subset $A$ of $M$ is said to be convex if $a,b\in A$ and $c\in M$ with $a<c<b$ then $c\in A$.
@@
 We say that $\mathcal{M}$ is o-minimal (weakly o-minimal) if every definable subset of $M$ is a finite union of intervals (convex sets), respectively.
-In o-minimal structures, Wilkie ([9]) proved every bounded definable open set is a finite union of open cells.
+Weakly o-minimality was introduced by Dickmann in [3].
+Wilkie ([12]) proved that in o-minimal expansions of a real closed field every bounded definable open set is a finite union of open cells.
 Andrews ([1]) proved that if an o-minimal structure admits CE-cell decomposition then any definable open set is expressed as a finite union of definable open cells.
-Edmundo et al ([3]) proved in semi-bounded o-minimal expansion of an ordered group, any definable open set is covered a finite union of open cells.
-The relation between strong cell decomposition in weakly o-minimal structures and open cell property in o-minimal structures is proved by [4].
+Edmundo et al ([6]) proved in semi-bounded o-minimal expansion of an ordered group, any definable open set is covered by a finite union of open cells.
+The relation between strong cell decomposition in weakly o-minimal structures and the open cell property in o-minimal structures is proved by [7].
 We consider a generalization of the above results in weakly o-minimal structures.
+Readers are expected to be familiar with fundamental results of o-minimality and weak o-minimality; see, for example, [2], [4], [9], [10].
 2. Weakly o-minimal expansions of real closed fields
-Let $\mathcal{M}=(M,<,\ldots)$ be an weakly o-minimal structure.
+Let $\mathcal{M}=(M,<,\ldots)$ be a weakly o-minimal expansion of a dense linear order without endpoints.
 For any subsets $C,D$ of $M$, we write $C<D$ if $c<d$ whenever $c\in C$ and $d\in D$.
 A pair $\langle C,D\rangle$ of non-empty subsets of $M$ is called a cut in $\mathcal{M}$ if $C<D,C\cup D=M$ and $D$ has no lowest element.
-A cut $\langle C,D\rangle$ is said to be definable in $\mathcal{M}$ if the sets $C,D$ are definable in $\mathcal{M}$.
+A cut $\langle C,D\rangle$ is said to be definable in $\mathcal{M}$ if the sets $C$ and $D$ are definable in $\mathcal{M}$.
 The set of all cuts definable in $\mathcal{M}$ will be denoted by $\overline{M}$.
@@
 Then we may treat $(M,<)$ as a substructure of $(\overline{M},<)$ by identifying an element $a\in M$ with the definable cut $\langle(-\infty,a],(a,+\infty)\rangle$.
-Morever, if $\mathcal{M}=(M,<,+,\ldots)$ is an expansion of an ordered group $(M,<,+)$, then the cut $\langle C,D\rangle$ is said to be valuational if there exists $\varepsilon>0$ such that for all $x\in C$ and $y\in D$, we have $y-x>\varepsilon$.
+Moreover, if $\mathcal{M}=(M,<,+,\ldots)$ is an expansion of an ordered group $(M,<,+)$, then the cut $\langle C,D\rangle$ is said to be valuational if there exists $\varepsilon>0$ such that for all $x\in C$ and $y\in D$, we have $y-x>\varepsilon$.
 Otherwise the cut $\langle C,D\rangle$ is said to be non-valuational.
@@
 A function $f:A\to\overline{M}\cup\{-\infty,+\infty\}$ is said to be definable if $f$ is a definable function from $A$ to $\overline{M}$, $f(x)=$-$\infty$ for all $x\in A$, or $f(x)=+\infty$ for all $x\in A$.
-A definable function $f:A\to\overline{M}\cup\{-\infty,+\infty\}$ is said to be regular if it takes all its values in one of the sets $M$, $\overline{M}-M$, $\{-\infty\}$, or $\{+\infty\}$.
-Definition 2.1 ([7]).
+We recall the notion of strong cells from [10].
+Definition 2.1.
 For every $m\in\mathbb{N}_{+}$, we define, by induction, strong cells $C\subseteq M^{m}$ and their completions $\overline{C}\subseteq\overline{M}^{m}$.
@@
   4. (4)
-    If $C\subseteq M^{m}$ is a $k$-strong cell and $f,g:C\to M\cup\{\pm\infty\}$ is continuous definable functions such that $f,g$ have continuous extensions $\overline{f},\overline{g}:\overline{C}\to\overline{M}$, where $\overline{f}(\overline{a})<\overline{g}(\overline{a})$ for any $\overline{a}\in\overline{C}$, then the set $(f,g)_{C}:=\{\langle\overline{a},b\rangle\in C\times M:f(\overline{a})<b<g(\overline{a})\}$ is called a $(k+1)$-strong cell in $M^{m+1}$.
+    If $C\subseteq M^{m}$ is a $k$-strong cell and $f,g:C\to\overline{M}\cup\{\pm\infty\}$ is continuous definable functions such that $f,g$ have continuous extensions $\overline{f},\overline{g}:\overline{C}\to\overline{M}\cup\{\pm\infty\}$, where $\overline{f}(\overline{a})<\overline{g}(\overline{a})$ for any $\overline{a}\in\overline{C}$, then the set $(f,g)_{C}:=\{\langle\overline{a},b\rangle\in C\times M:f(\overline{a})<b<g(\overline{a})\}$ is called a $(k+1)$-strong cell in $M^{m+1}$.
 The completion of $(f,g)_{C}$ in $\overline{M}^{m+1}$ is defined as $(\overline{f},\overline{g})_{\overline{C}}:=\{\langle\overline{a},b\rangle\in\overline{C}\times\overline{M}:\overline{f}(\overline{a})<b<\overline{g}(\overline{a})\}$.
-A strong cell $C\subseteq M^{m}$, $m\geq 2$, is called a refined strong cell if each of the boundary functions appearing in its definition is regular.
+  5. (5)
+    Every $k$-strong cell is of the form given in (1) through (4).
+We say that $C\subseteq M^{m}$ is a strong cell in $M^{m}$ if there exists a non-negative integer $k$ such that $C$ is a $k$-strong cell in $M^{m}$.
+A function $f:C\to\overline{M}$ is said to be a cell-defining function if $\Gamma(f)_{C}$ is a strong cell or there is a definable function $g:C\to\overline{M}$ so that $(f,g)_{C}$ or $(g,h)_{C}$ is a strong cell.
+A strong cell $C\subseteq M^{m}$, $m\geq 2$, is called a refined strong cell if each of the cell-defining functions appearing in its definition assumes all its values in one of the following sets: $M$, $\overline{M}\setminus M$, $\{-\infty\}$, or $\{+\infty\}$.
 Refined strong cells in $M$ coincide with strong cells in $M$.
+Note that the definitions of strong cells in [10] and [11] are not identical.
+A strong cell in [11] is called a refined strong cell in [10].
+In this paper, we follow the terminology of [10].
 Let $C\subseteq M^{m}$ be a strong cell and $f:C\to\overline{M}$ definable.
@@
 Similarly, we can define the refined strong cell decomposition property.
-The following result follows from Lemma 2.6 and Corollary 2.16 of [7].
+The following result follows from Lemma 2.6, Corollary 2.16, and Corollary 2.17 of [10].
 Fact 2.3.
-Let $\mathcal{M}=(M,<,+,\ldots)$ be a weakly o-minimal expansion of an ordered group. Then the structure $\mathcal{M}$ is a non-valuational if and only if it has a strong cell decomposition property.
-Let $\mathcal{C}$ and $\mathcal{D}$ are refined strong cell decompositions of $M^{m}$.
-We say that $\mathcal{D}$ refines $\mathcal{C}$ (denoted by $\mathcal{C}\prec\mathcal{D}$) if every cell in $\mathcal{C}$ is a union of some cells in $\mathcal{D}$
-The following statement is Fact 2.1 of [8].
-Fact 2.4.
-Let $k,m\in\mathbb{N}_{+}$.
-If $X_{1},\ldots,X_{k}\subseteq M^{m}$ are definable sets, then there exists a smallest refined strong cell decomposition $\mathcal{D}$ of $M^{m}$ partitioning each of $X_{1},\ldots,X_{k}$.
+Let $\mathcal{M}=(M,<,+,\ldots)$ be a weakly o-minimal expansion of an ordered group.
+Then the structure $\mathcal{M}$ is non-valuational if and only if it has the refined strong cell decomposition property.
+The following definition appears in [11].
+Definition 2.4.
+Let $\mathcal{M}=(M,<,\ldots)$ be a weakly o-minimal structure with the refined strong cell decomposition property.
+For any $m\in\mathbb{N}_{+}$ and $i_{1},\dots,i_{m}\in\{0,1\}$, we define basic $\langle i_{1},\dots,i_{m}\rangle$-cells in $\overline{M}^{m}$ inductively.
+If $1\leq j_{1}<\dots<j_{k}\leq m$, let
+$\rho^{m}_{j_{1},\dots,j_{k}}:\overline{M}^{m}\to\overline{M}^{k}$
+denote the projection onto the coordinates $j_{1},\dots,j_{k}$. When $m$ is clear from the context, we simply write $\rho_{j_{1},\dots,j_{k}}$.
+  1. (1)
+    A singleton of $\overline{M}^{m}$ is called a basic $\langle 0,\dots,0\rangle$-cell in $\overline{M}^{m}$.
+  2. (2)
+    If $C\subseteq M$ is a strong $\langle 1\rangle$-cell, then $\overline{C}$ is called a basic $\langle 1\rangle$-cell. Note that
+    $\rho_{1}(\overline{C})\cap M=\overline{C}\cap M=C$
+    is an open strong cell in $M$.
+  3. (3)
+    If $C=\{\overline{a}\}\subseteq\overline{M}^{m}$ and $I$ is a basic $\langle 1\rangle$-cell in $\overline{M}$, then $C\times I$ is called a basic $\langle 0,\dots,0,1\rangle$-cell in $\overline{M}^{m+1}$.
+Note that
+    $\rho_{m+1}(C\times I)\cap M=I\cap M$
+    is an open strong cell in $M$.
+  4. Assume that $i_{1},\dots,i_{m}\in\{0,1\}$ with $i_{1}+\dots+i_{m}>0$, and suppose that basic $\langle i_{1},\dots,i_{m}\rangle$-cells in $\overline{M}^{m}$ have already been defined. Let
+    $\{j_{1},\dots,j_{k}\}=\{j\in\{1,\dots,m\}:i_{j}=1\},\qquad j_{1}<\dots<j_{k},$
+    and suppose that for every basic $\langle i_{1},\dots,i_{m}\rangle$-cell $C\subseteq\overline{M}^{m}$, the set
+    $\rho_{j_{1},\dots,j_{k}}(C)\cap M^{k}$
+    is an open strong cell in $M^{k}$.
+  5. (4)
+    Let $C\subseteq\overline{M}^{m}$ be a basic $\langle i_{1},\dots,i_{m}\rangle$-cell, and set
+    $D:=\rho_{j_{1},\dots,j_{k}}(C)\cap M^{k},$
+    which is an open strong cell in $M^{k}$. If $f$ is a strongly continuous definable function from $D$ to $M$ or from $D$ to $\overline{M}\setminus M$, then
+    $\Gamma\bigl(\overline{f}\circ\rho_{j_{1},\dots,j_{k}}\!\restriction_{C}\bigr)$
+    is called a basic $\langle i_{1},\dots,i_{m},0\rangle$-cell in $\overline{M}^{m+1}$. Note that
+    $\rho_{j_{1},\dots,j_{k}}\!\left(\Gamma\bigl(\overline{f}\circ\rho_{j_{1},\dots,j_{k}}\!\restriction_{C}\bigr)\right)\cap M^{k}=D$
+    is an open strong cell in $M^{k}$.
+  6. (5)
+    Let $C\subseteq\overline{M}^{m}$ be a basic $\langle i_{1},\dots,i_{m}\rangle$-cell, and set
+    $D:=\rho_{j_{1},\dots,j_{k}}(C)\cap M^{k},$
+    an open strong cell in $M^{k}$. If $f,g:D\to\overline{M}\cup\{\pm\infty\}$ are strongly continuous definable functions such that all values of $f$ and $g$ lie in one of the sets
+    $\{-\infty\},\quad M,\quad\overline{M}\setminus M,\quad\{+\infty\},$
+    and
+    $\overline{f}(\overline{x})<\overline{g}(\overline{x})\quad\text{for all }\overline{x}\in\overline{D},$
+    then the set
+    $(\overline{f}\circ\rho_{j_{1},\dots,j_{k}},\overline{g}\circ\rho_{j_{1},\dots,j_{k}})_{C}:=\left\{\langle\overline{a},b\rangle\in C\times\overline{M}:(\overline{f}\circ\rho_{j_{1},\dots,j_{k}})(\overline{a})<b<(\overline{g}\circ\rho_{j_{1},\dots,j_{k}})(\overline{a})\right\}$
+    is called a basic $\langle i_{1},\dots,i_{m},1\rangle$-cell in $\overline{M}^{m+1}$. Note that
+    $\rho_{j_{1},\dots,j_{k}}\left[(\overline{f}\circ\rho_{j_{1},\dots,j_{k}},\overline{g}\circ\rho_{j_{1},\dots,j_{k}})_{C}\right]\cap M^{k+1}=(f,g)_{D}$
+    is an open strong cell in $M^{k+1}$.
+In the standard way, we define a basic cell decomposition of $\overline{M}^{m}$ as a finite partition of $\overline{M}^{m}$ into basic cells.
 For any $n\in\mathbb{N}_{+}$ and any refined strong cell $C\subseteq M^{n}$, we denote by $R_{C}$ an $n$-ary relational symbol.
@@
 If additionally $Y\subseteq M^{m}$ is definable in $\mathcal{M}$, then $X\cap Y$ is definable in $\mathcal{M}$.
-The following definition is Definition 3.1 of [6].
-Definition 2.5.
-Let $m\in\mathbb{N}_{+}$.
-Suppose that $X$ is a defianble subset of $M^{m}$ and $\mathcal{C}$ is the smallest refined strong cell decomposition of $M^{m}$ partitioning $X$.
-We define the completion of $X$ in $\overline{M}^{m}$ as $\overline{X}:=\bigcup\{\overline{C}:C\in\mathcal{C}\wedge C\subseteq X\}$.
-Remark 2.6.
-Note that there exist sets $C,D$ such that $C\subseteq D$ but $\overline{C}\not\subseteq\overline{D}$.
-Let $R_{alg}$ denote the ordered field of real algebraic numbers and let the unary predicate symbol $P$ be interpreted by the convex set $S=(-\pi,\pi)\cap R_{alg}$.
-Then, by Proposition 2.1 from [5], the structure $\mathfrak{R}=(R_{alg},<,+,\cdot,P)$ is non-valuational weakly o-minimal.
+The following result follows from Proposition 2.3 of [11].
+Fact 2.5.
+Assume that $\mathcal{M}=(M,<,\dots)$ is a weakly o-minimal structure with the strong cell decomposition property and $\overline{\mathcal{M}}=(\overline{M},<,\dots)$ is its canonical o-minimal extension.
+Let $m,k\in\mathbb{N}_{+}$.
+If $X_{1},\dots,X_{k}\subseteq\overline{M}^{m}$ are sets definable in $\overline{{\mathcal{M}}}$, then there exists a basic cell decomposition of $\overline{M}^{m}$ into finitely many basic cells partitioning each of the sets $X_{1},\ldots,X_{k}$.
+Proposition 2.6.
+Suppose that $\mathcal{M}=(M,<,\ldots)$ is a weakly o-minimal structure with the refined strong cell decomposition property.
+Let $n\in\mathbb{N}_{+}$.
+Suppose that $X$ is a definable open subset of $M^{n}$.
+Then there exists an $\overline{{\mathcal{M}}}$-definable open set $Y\subseteq\overline{M}^{n}$ such that $X=Y\cap M^{n}$.
+Proof.
+Take a refined strong cell decomposition of $X$, so that
+$X=C_{1}\cup\dots\cup C_{k}.$
+By Fact 2.5, there exists a basic cell decomposition $\mathcal{C}$ of $\overline{M}^{n}$ partitioning each of the sets $\overline{C_{1}},\dots,\overline{C_{k}}$. Define
+$\hat{X}=\bigcup\left\{D\in\mathcal{C}:\exists i\ (D\subseteq\overline{C_{i}})\right\}\;\cup\;\bigcup\left\{D\in\mathcal{C}:\bigl(\forall i\ (D\cap\overline{C_{i}}=\emptyset)\bigr)\wedge(D\cap M^{n}=\emptyset)\right\}.$
+[…]
+By the definition of $\hat{X}$, the inclusion $X\supseteq Y\cap M^{n}$ is clear.
+[…]
+If $C\subseteq\overline{C_{i}}$ for some $i$, then $B\subseteq\overline{C_{i}}\subseteq\hat{X}$, a contradiction.
+Thus for every $i$ we have $C\cap\overline{C_{i}}=\emptyset$, and hence $C\cap X=\emptyset$.
+First, consider the case where $C$ is a basic $\langle 0,\dots,0,1\rangle$-cell. Then there exists an open interval $I\subseteq\overline{M}$ such that
+$C=(a_{1},\dots,a_{n-1})\times I.$
+Since $a\in X\subseteq M^{n}$ and $a\in\operatorname{cl}_{\overline{M}}(C)$, every open box $U$ in $M^{n}$ containing $a$ satisfies $C\cap U\neq\emptyset$, equivalently $(C\cap M^{n})\cap U\neq\emptyset$.
+Because $C\cap X=\emptyset$, we have $(C\cap M^{n})\cap X=\emptyset$.
+Hence $a\notin\operatorname{int}_{M}(X)$, contradicting that $X$ is open in $M^{n}$.
+Next, consider the case where $C$ is a basic $\langle i_{1},\dots,i_{n-1},0\rangle$-cell and $i_{1}+\dots+i_{n-1}>0$.
+That is, there exists a basic $\langle i_{1},\dots,i_{n-1}\rangle$-cell $\hat{D}\subseteq\overline{M}^{n-1}$, and if we set
+$\{j_{1},\ldots,j_{k}\}=\{\,j\in\{1,\dots,n-1\}:i_{j}=1\,\},\qquad j_{1}<\dots<j_{k},$
+then
+$D=\rho_{j_{1},\ldots,j_{k}}(\hat{D})\cap M^{k}$
+is an open strong cell in $M^{k}$. Moreover, there exists a strongly continuous definable function $f$ from $D$ to $M$ or from $D$ to $\overline{M}\setminus M$ such that
+[…]
+Assume that $f:D\to M$.
+Recall that $C\cap\overline{C_{i}}=\emptyset$ for every $1\leq i\leq k$.
+If $C\cap M^{n}=\emptyset$, $C\subseteq\hat{X}$ by the definition of $\hat{X}$.
+Hence $B\subseteq C\subseteq\hat{X}$, a contradiction.
+Assume that $C\cap M^{n}\neq\emptyset$.
+In the construction of a function-type cell in the basic cells, all cell-defining functions take values in $M$.
+We claim that $a\in\operatorname{cl}_{\overline{M}}(C\cap M^{n})$.
+Let $U\subseteq M^{n}$ be an arbitrary open box containing $a$.
+Since $a\in\operatorname{cl}_{\overline{M}}(C)$, there exists a point $x=(x_{1},\dots,x_{n})\in\overline{U}\cap C$.
+Set $\tilde{C}=\rho_{j_{1},\ldots,j_{k}}(C)$ and $V=\rho_{j_{1},\ldots,j_{k}}(U)$.
+Then $\tilde{C}$ is open in $\overline{M}^{k}$, and $V$ is an open box in $M^{k}$.
+Let $\tilde{x}=(x_{j_{1}},\dots,x_{j_{k}})$.
+Then $\tilde{x}\in\overline{V}\cap\tilde{C}$.
+Since $\tilde{C}$ is open in $\overline{M}^{k}$, there exists an open box $W$ in $M^{k}$ such that $\tilde{x}\in\overline{W}\subseteq\overline{V}\cap\tilde{C}$.
+Choose $\tilde{y}\in W$.
+Since $f:D\to M$, there exists $y\in(C\cap M^{n})\cap U$ such that $\rho_{j_{1},\ldots,j_{k}}(y)=\tilde{y}$.
+Therefore $(C\cap M^{n})\cap U\neq\emptyset$, and hence $a\in\operatorname{cl}_{\overline{M}}(C\cap M^{n})$.
+However, since $C\cap X=\emptyset$, we have $(C\cap M^{n})\cap X=\emptyset$, and hence $a\notin\operatorname{int}_{M}(X)$.
+This contradicts the assumption that $X$ is open in $M^{n}$.
+Assume instead that $f:D\to\overline{M}\setminus M$. Then $C\cap M^{n}=\emptyset$, and similarly $C\subseteq\hat{X}$. Hence $B\subseteq C\subseteq\hat{X}$, a contradiction.
+Therefore the function-type case leads to a contradiction.
+Finally, consider the case where $C$ is a basic $\langle i_{1},\dots,i_{n-1},1\rangle$-cell and $i_{1}+\dots+i_{n-1}>0$.
+As before, there exist a basic $\langle i_{1},\dots,i_{n-1}\rangle$-cell $\hat{D}\subseteq\overline{M}^{n-1}$ and an open strong cell $D\subseteq M^{k}$.
+Moreover, there exist strongly continuous functions $f,g:D\to\overline{M}\cup\{\pm\infty\}$ such that all values of $f$ and $g$ lie in exactly one of $\{-\infty\},\ M,\ \overline{M}\setminus M,\ \{\infty\}$, and $\overline{f}(x)<\overline{g}(x)$ for any $x\in\overline{D}$, and
+$C=(\overline{f}\circ\rho_{j_{1},\ldots,j_{k}},\ \overline{g}\circ\rho_{j_{1},\ldots,j_{k}})_{\hat{D}}$.
+Recall that $C\cap\overline{C_{i}}=\emptyset$ for every $1\leq i\leq k$.
+If $C\cap M^{n}=\emptyset$, $C\subseteq\hat{X}$ by the definition of $\hat{X}$.
+Hence $B\subseteq C\subseteq\hat{X}$, a contradiction.
+Assume $C\cap M^{n}\neq\emptyset$.
+As in the previous case, we have $a\in\operatorname{cl}_{\overline{M}}(C\cap M^{n})$.
+Hence every open box $U$ in $M^{n}$ containing $a$ meets $C\cap M^{n}$.
+Because $C\cap X=\emptyset$, we have $(C\cap M^{n})\cap X=\emptyset$, so $a\notin\operatorname{int}_{M}(X)$, contradicting openness of $X$ in $M^{n}$.
+This completes the proof.
+∎
+We recall the notion of open cell property from [1].
+An o-minimal structure $\mathcal{M}=(M,<,\dots)$ is said to have the open cell property if every non-empty definable open subset of $M^{n}$ is a union of finitely many open cells.
+A weakly o-minimal structure $\mathcal{M}=(M,<,\dots)$ is said to have the open cell property if every non-empty definable open subset of $M^{n}$ is a union of finitely many open strong cells.
+The following example shows that the open cell property fails for refined strong cells.
+Example 2.7.
+Let $R_{\mathrm{alg}}$ denote the ordered field of real algebraic numbers and let the unary predicate symbol $P$ be interpreted by the convex set $S=(-\pi,\pi)\cap R_{\mathrm{alg}}$.
+Then, by Proposition 2.1 of [9], the structure $\mathfrak{R}=(R_{\mathrm{alg}},<,+,\cdot,P)$ is non-valuational weakly o-minimal.
 So the structure $\mathfrak{R}$ has the refined strong cell decomposition property.
-Consider the functions $f,g:I=(0,4)\to\mathfrak{R}$ defined by $f(x)=x$ and $g(x)=x+1$, and the function $h:J=(0,\pi)\to\mathfrak{R}$ defined by $h(x)=x+2$.
-Let $C=\Gamma(f)_{I}\cup\Gamma(g)_{I}$ and $D=\Gamma(f)_{I}\cup\Gamma(g)_{I}\cup\Gamma(h)_{J}$.
-Then, we have $C\subseteq D$ but $\overline{C}\not\subseteq\overline{D}$.
-Remark 2.7.
-Let $m\in\mathbb{N}_{+}$.
-Suppose that $X$ is a defianble subset of $M^{m}$.
-Let $\mathcal{C}$ be the smallest refined strong cell decomposition of $M^{m}$ partitioning $X$ such that $X=\bigcup_{i=1}^{k}C_{i}$ and $\overline{X}=\bigcup_{i=1}^{k}\overline{C_{i}}$ with $C_{1},\ldots,C_{k}\in\mathcal{C}$.
-If $X$ is open in $M$, then $\overline{X}$ is open in $\overline{M}$.
-Proposition 2.8 (Weak definable choice).
-Suppose that $\mathcal{M}=(M,<,+,\ldots)$ is a weakly o-minimal expansion of an ordered group $(M,<,+)$.
-Let $S\subseteq M^{m+n}$ be definable and $\pi:M^{m+n}\to M^{m}$ the projection of the first $m$ coordinates.
-Then, there exists a definable map $f:\pi(S)\to\overline{M}^{n}$ such that $\Gamma(f)\subseteq\overline{S}$.
-Proof.
-Let $\mathcal{C}$ be the smallest refined strong cell decomposition of $M^{m+n}$ partitioning $S$ such that $S=\bigcup_{i=1}^{k}C_{i}$ and $\overline{S}=\bigcup_{i=1}^{k}\overline{C_{i}}$ with $C_{1},\ldots,C_{k}\in\mathcal{C}$.
-By the o-minimality of $\overline{{\mathcal{M}}}$ and Proposition 6.1.2 of [2], for any $C\in\mathcal{C}$ with $C\subseteq S$, there exists a map $\overline{f}_{C}:\overline{\pi(C)}\to\overline{M}^{n}$ such that $\Gamma(\overline{f}_{C})\subseteq\overline{C}$ and $\overline{f}_{C}$ is definable in $\overline{{\mathcal{M}}}$.
-By Propisition 2.3 from [8], $\overline{f}_{C}|_{\overline{\pi(C)}\cap M^{m}}:\overline{\pi(C)}\cap R^{m}\to\overline{M}^{n}$ is definable in $\mathcal{M}$}.
-By $S=\bigcup_{i=1}^{k}C_{i}$ and $\overline{S}=\bigcup_{i=1}^{k}\overline{C_{i}}$, this completes the proof.
-∎
-Proposition 2.9.
-Suppose that $\mathcal{M}=(M,<,+,\ldots)$ is a weakly o-minimal expansion of an ordered group $(M,<,+)$.
-Let $S\subseteq M^{m+n}$ be definable and $\overline{\pi}:\overline{M}^{m+n}\to\overline{M}^{m}$ the projection of the first $m$ coordinates.
-Then, there exists a map $\overline{g}:\overline{\pi}(\overline{S})\to\overline{S}$ such that $\overline{\pi}\circ\overline{g}=\mathrm{id}_{\overline{\pi}(\overline{S})}$ and $\overline{g}$ is definable in $\overline{{\mathcal{M}}}$.
-Proof.
-Let $\pi:=\overline{\pi}|_{M^{m+n}}$.
-Suppose that $\mathcal{C}$ is the smallest refined strong cell decomposition of $M^{m+n}$ partitioning $S$ such that $S=\bigcup_{i=1}^{k}C_{i}$ and $\overline{S}=\bigcup_{i=1}^{k}\overline{C_{i}}$ with $C_{1},\ldots,C_{k}\in\mathcal{C}$.
-By the o-minimality of $\overline{{\mathcal{M}}}$ and Proposition 6.1.2 of [2], for any $C\in\mathcal{C}$ with $C\subseteq S$, there exists a map $\overline{f}_{C}:\overline{\pi(C)}\to\overline{M}^{n}$ such that $\Gamma(\overline{f}_{C})\subseteq\overline{C}$ and $\overline{f}_{C}$ is definable in $\overline{{\mathcal{M}}}$.
-Let $\overline{f}=\bigcup_{i=1}^{k}\overline{f}_{C_{i}}$.
-Then we have $\overline{f}:\overline{\pi}(\overline{S})\to\overline{M}^{n}$ and $\Gamma(\overline{f})\subseteq\overline{S}$.
-We define the $\overline{\mathcal{M}}$-definable map $\overline{g}:\overline{\pi}(\overline{S})\to\overline{S}$ by $\overline{g}(x)=(x,\overline{f}(x))$.
-Then it follows $\overline{\pi}\circ\overline{g}=\mathrm{id}_{\overline{\pi}(\overline{S})}$.
-∎
-Example 2.10.
-Let $R_{alg}$ denote the ordered field of real algebraic numbers and let the unary predicate symbol $P$ be interpreted by the convex set $S=(-\pi,\pi)\cap R_{alg}$.
-Then, by Proposition 2.1 from [5], the structure $\mathfrak{R}=(R_{alg},<,+,\cdot,P)$ is non-valuational weakly o-minimal.
-So the structure $\mathfrak{R}$ has the refined strong cell decomposition property.
-Consider the functions $f,g:I=(-1,1)\to\overline{\mathfrak{R}}$ defined by f(x)=4f(x)=-4 and $g(x)=\pi x$.
-Then, the set $(f,g)_{I}$ is an open strong cell but not an open refined strong cell.
+Consider the functions $f,g:I=(-1,1)\to\overline{R}_{\mathrm{alg}}$ defined by f(x)=4f(x)=-4 and $g(x)=\pi x$.
+The set $(f,g)_{I}$ is an open strong cell but not a refined strong cell because $g(0)=0\in R_{\mathrm{alg}}$ and $g(x)\notin R_{\mathrm{alg}}$ for $x\neq 0$.
 Moreover, $(f,g)_{I}$ cannot be expressed as a finite union of open refined strong cells.
-Lemma 2.11 ([5]).
-Let $\mathcal{M}$ be a weakly o-minimal expansion of an ordered field. Then $\mathcal{M}$ is a real closed field.
-Fact 2.12 ([9]).
+Fact 2.8 ([12, Theorem 1.3]).
 Suppose that $\mathcal{M}=(M,<,+,\cdot,\ldots)$ is an o-minimal expansion of a real closed field.
@@
 Proof.
-Let $\mathcal{C}$ be the smallest refined strong cell decomposition of $M^{n}$ partitioning $U$ such that $U=\bigcup_{i=1}^{k}C_{i}$ and $\overline{U}=\bigcup_{i=1}^{k}\overline{C_{i}}$ with $C_{1},\ldots,C_{k}\in\mathcal{C}$.
-By Remark 2.7, the set $\overline{U}$ is a definable bounded open subset of $\overline{M}^{n}$.
-Then, by Fact 2.12, there exist open cells $D_{1},\dots,D_{l}$ such that $\overline{U}=D_{1}\cup\dots\cup D_{l}$.
+By Fact 2.3, the structure $\mathcal{M}$ has the refined strong cell decomposition property.
+Since $U$ is a definable bounded open subset of $M^{n}$, by Proposition 2.6, there exists an $\overline{{\mathcal{M}}}$-definable bounded open set $V\subseteq\overline{M}^{n}$ such that $U=V\cap M^{n}$.
+By Fact 2.8, there exist open cells $D_{1},\dots,D_{l}$ such that $V=D_{1}\cup\dots\cup D_{l}$.
+Because each $D_{i}\cap M^{n}$ is an open strong cell and ${U}=(D_{1}\cap M^{n})\cup\dots\cup(D_{l}\cap M^{n})$, this completes the proof.
+∎
+3. Weakly o-minimal expansions of ordered groups
+Let $\mathcal{M}=(M,<,+,\cdots)$ be a non-valuational weakly o-minimal expansion of an ordered group.
+If $\mathcal{M}$} is o-minimal, we say that $\mathcal{M}$ is semi-bounded if there exists no definable bijection between a bounded interval and an unbounded interval [5].
+Theorem 3.1 ([6]).
+Let $\mathcal{M}=(M,<,+,\cdots)$ be a semi-bounded o-minimal expansion of an ordered group. Then every non-empty definable open set is a finite union of open cells.
+Theorem 3.2.
+Consider a non-valuational weakly o-minimal expansion $\mathcal{M}=(M,<,+,\dots)$ of an ordered group.
+Let $U$ be a definable open subset of $M^{n}$ and assume that $\overline{{\mathcal{M}}}$ is semi-bounded.
+Then there exist finitely many open strong cells whose union is $U$.
+Proof.
+By Proposition 2.6, there exists an $\overline{{\mathcal{M}}}$-definable open subset $V$ of $\overline{M}^{n}$ such that $U=V\cap M^{n}$.
+By Theorem 3.1, there exist open cells $D_{1},\dots,D_{l}$ such that $V=D_{1}\cup\dots\cup D_{l}$.
 Since each $D_{i}\cap M^{n}$ is an open strong cell and ${U}=(D_{1}\cap M^{n})\cup\dots\cup(D_{l}\cap M^{n})$, this completes the proof.
 ∎
-3. Weakly o-minimal expansions of ordered groups
-Let $\mathcal{M}$ be a non-vaulational weakly o-minimal expansion $\mathcal{M}=(M,<,+,0,\cdots)$ of an ordered group.
-We say that $\mathcal{M}$ is semisemi-boundedbounded if there exists no definable bijection between a bounded interval and a unbounded interval.
-Theorem 3.1 ([3]).
-Let $\mathcal{M}$ be a semi-bounded o-minimal expansion $\mathcal{M}=(M,<,+,0,\cdots)$ of an ordered group. Then every non-empty definable open set is a finite union of open cells.
+We obtained an improvement of Theorem 3.2 inspired by a private communication with Fujita [8].
 Theorem 3.3.
-Consider a semi-bounded non-valuational weakly o-minimal expansion $\mathcal{M}=(M,<,+,\dots)$ of an ordered group.
-Let $U$ be a definable open subset of $M^{n}$.
-Then there exist finitely many strong open cells whose union is $U$.
+Let $\mathcal{M}=(M,<,+,\dots)$ be a non-valuational weakly o-minimal expansion of an ordered group. Then the following three conditions are equivalent.
+(1) For any bounded definable set $I\subseteq M$ and for every $\mathcal{M}$-definable function $f:I\to\overline{M}$, $f(I)$ is bounded.
+(2) The canonical o-minimal extension $\overline{\mathcal{M}}$ of $\mathcal{M}$ is semi-bounded.
+(3) Any $\mathcal{M}$-definable open set is a finite union of open strong cells.
 Proof.
-Take the smallest refined strong cell decomposition $\mathcal{C}$ of $M^{n}$ partitioning $U$ such that $U=\bigcup_{i=1}^{k}C_{i}$ and $\overline{U}=\bigcup_{i=1}^{k}\overline{C_{i}}$ with $C_{1},\ldots,C_{k}\in\mathcal{C}$.
-By Remark 2.7, the set $\overline{U}$ is a definable open subset of $\overline{M}^{n}$.
-Then, by Theorem 3.1, there exist open cells $D_{1},\dots,D_{l}$ such that $\overline{U}=D_{1}\cup\dots\cup D_{l}$.
-Since each $D_{i}\cap M^{n}$ is an open strong cell and ${U}=(D_{1}\cap M^{n})\cup\dots\cup(D_{l}\cap M^{n})$, this completes the proof.
+$(1)\Rightarrow(2)$ Assume that (2) does not hold.
+Then we have a bounded open interval $J_{1}$, an unbounded interval $J_{2}$ and an $\overline{{\mathcal{M}}}$-definable bijection $g:J_{1}\to J_{2}$.
+Since $\overline{{\mathcal{M}}}$ is o-minimal and by the monotonicity theorem, we may assume $g$ is continuous and strictly monotone.
+Set $I=J_{1}\cap M$, $f=g|I$.
+Then $f$ is $\mathcal{M}$-definable.
+Since $M$ is dense in $\overline{M}$, $f(I)$ is unbounded.
+This contradicts (1).
+$(2)\Rightarrow(1)$ We show the contradiction.
+Suppose that $I$ is a bounded subset of $M$ and $f:I\to\overline{M}$ is a function definable in $\mathcal{M}$.
+Assume that $f(I)$ is unbounded.
+By Proposition 3.1 of [11], we can decompose $I$ into $I=X\cup J_{1}\cup\dots\cup J_{k}$ such that $X$ is finite, each $J_{i}$ is open convex set and $f|J_{i}$ is monotone and strongly continuous.
+We have the fact that some $f(J_{i})$ is unbounded.
+Replacing $f$ by $f|J_{i}$, we can assume that $f$ is monotone and strongly continuous.
+Then the unique extension $\overline{f}:\overline{I}\to\overline{M}$ is monotone and continuous.
+This is an $\overline{{\mathcal{M}}}$-definable bijection between a bounded open interval $\overline{I}$ and an unbounded open interval $\overline{f}(\overline{I})$.
+This is a contradiction.
+$(2)\Rightarrow(3)$ By Theorem 3.2, this implication follows.
+$(3)\Rightarrow(2)$ Assume that (2) does not hold.
+Then we have an $\overline{{\mathcal{M}}}$-definable bijection $f:J_{1}\to J_{2}$ between a bounded open interval $J_{1}$ and an unbounded interval $J_{2}$.
+Since $\overline{{\mathcal{M}}}$ is o-minimal and by the monotonicity theorem, we may assume that $f$ is strictly monotone and continuous.
+We may further assume that $J_{2}$ is bounded below and $f$ is strictly decreasing.
+By shifting, we can set $J_{1}=(0,b)$.
+Let $Y$ be the $\overline{{\mathcal{M}}}$-definable open set
+$Y=\{(x,y)\in\overline{M}^{2}:(-b<x<b)\land((x<0)\to(y<f(-x)))\land((x>0)\to(y<f(x)))\}$.
+Set $X=Y\cap M^{n}$. Then $X$ is $\mathcal{M}$-definable. We prove that $X$ is not a finite union of open strong cells.
+Assume $X$ is a finite union of open strong cells.
+Note that $\{0\}\times M$ is contained in $X$.
+Thus there exists an open strong cell $C$ and $r\in M$ such that $\{0\}\times(r,\infty)\subseteq C\subseteq X$.
+Since $\{0\}\times(r,\infty)\subseteq C$, we can find a definable open convex set $I$ with $0\in I$ and a strongly continuous definable function $g:I\to\overline{M}\cup\{-\infty\}$ such that $C=\{(x,y)\in M^{2}:x\in I,y>g(x)\}$.
+Hence $C$ contains a point $(x_{0},y_{0})$ with $0<x_{0}<b$ and $y_{0}>f(x_{0})$.
+It contradicts $C\subseteq X$.
 ∎